OpenStudy (anonymous):

Solve the following equation. x^2-3x=0

OpenStudy (anonymous):

take x as common term

OpenStudy (anonymous):

x^2-3x = 0 x(x-3) = 0 x=0, x=3

OpenStudy (anonymous):

x^2 - 3x = 0 x (x-3) = 0 now u can find the solutions..

OpenStudy (anonymous):

ok thanx